OpenStudy (anonymous):

Find the area of the shaded sector below. Use your calculator and round your final answer to the nearest hundredth.

OpenStudy (anonymous):

@Wendy.Ivette11714

OpenStudy (mathstudent55):

The central angle of a sector has the same measure as the arc's degree measure.

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Since the degree measure of the arc is 60 deg, the central angle of that sector is also 60 deg. Ok so far?

OpenStudy (mathstudent55):

The formula for the area of a sector is: \(\large A_{sector} = \dfrac{n}{360^o}\pi r^2\)

OpenStudy (jango_in_dtown):

area of the circle = pi 6^2 which is when the angle is 360 degree so the area for 60 degree is pi 6^2 *60/360= 6 pi=18.850

OpenStudy (mathstudent55):

n is the measure in degrees of the central angle of the sector.

OpenStudy (anonymous):

so confusing.....

OpenStudy (mathstudent55):

It's not confusing. Just follow the formula. You are given a central angle of 60 deg. That goes in for n in the formula. You are given a radius of 6 cm. That goes in for r in the formula. For pi use 3.14159. \(\large A_{sector} = \dfrac{n}{360^o}\pi r^2\) \(\large A_{sector} = \dfrac{60^o}{360^o}(3.14159)(6~cm)^2\)

OpenStudy (mathstudent55):

Remember to square the 6 first. Then multiply by 3.14159 (or use the pi key in your calculator if you have one) and multiply by 60. Finally divide by 360. What do you get?